OpenStudy (anonymous):

square root of 30 times square root of 10

OpenStudy (anonymous):

2 root 10

OpenStudy (anonymous):

how do i soulve it

OpenStudy (anonymous):

excuse me

OpenStudy (anonymous):

\[\sqrt{3(10)}\sqrt{10}\]\[=10\sqrt{3}\]